# RomanLens: Latent Romanization and its role in Multilinguality in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) exhibit remarkable multilingual generalization despite being predominantly trained on English-centric corpora. A fundamental question arises: how do LLMs achieve such robust multilingual capabilities? For non-Latin script languages, we investigate the role of romanization—the representation of non-Latin scripts using Latin characters—as a bridge in multilingual processing. Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in romanized form before transitioning to native script, a phenomenon we term Latent Romanization. Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and romanized scripts, suggesting a shared underlying representation. Additionally in translation towards non Latin languages, our findings reveal that when the target language is in romanized form, its representations emerge earlier in the model’s layers compared to native script. These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of romanization in facilitating language transfer. Our work provides new directions for potentially improving multilingual language modeling and interpretability.
RomanLens: Latent Romanization and its role in Multilinguality in LLMs
@@
Anoop Kunchukuttan1,3,6,
Mitesh M. Khapra1,3,
@@ 1 Introduction
In this study, we focus on languages written in non-Latin scripts. We hypothesize that LLMs leverage romanized forms of non-Latin script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations. Romanization—the representation of non-Latin scripts using Latin characters—may facilitate this process by aligning non-English languages more closely with English. Supporting this, Husain et al. (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations. We investigate whether LLMs indeed use romanization as a bridge between language-agnostic concepts and language-specific outputs given its potential implications for understanding multilingual processing in LLMs.
@@
To further probe this phenomenon, we employ activation patching Ghandeharioun et al. (2024); Variengien and Winsor (2023); Chen et al. (2024); Dumas et al. (2024), a technique that replaces activations from one forward pass with another to analyze the resulting outputs (see Section 3.2). Dumas et al. (2024) found that LLMs process language and conceptual information as distinct entities. Building on this, we perform layerwise activation patching between romanized and native-script inputs to examine whether LLMs encode conceptual information similarly across scripts.
, ’pa’ is its romanized representation). Note that push is a prefix of pushpam, the romanized form of the translation of “flower”; also, poo is another romanized translation of “flower.”
@@
1.
Latent Romanization Across Layers: During multilingual next-token generation, intermediate layers occasionally represent tokens in romanized form before resolving to native script (Figure 1). We term this phenomenon Latent Romanization.
2.
Consistent Semantic Encoding Across Scripts: Activation patching experiments reveal that LLMs encode semantic concepts similarly, regardless of whether the input is in native or romanized script.
3.
Earlier Emergence of Target Representations: When translating into romanized versus native script, romanized target representations emerge earlier in the model’s layers—typically one or two layers prior to native script representations (Figure 2). This suggests that romanization facilitates faster progression toward language-specific embeddings.
@@ 2 Related Work
Recent studies investigate latent representations in multilingual tasks, examining whether English emerges as a latent language in English-centric LLMs Wendler et al. (2024) and how the composition of training corpus mixtures influences latent representations Zhong et al. (2024). Recent investigations analyze how LLMs handle multilingual capabilitiesZhao et al. (2024b). Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al. (2023) and direct logit attribution Elhage et al. (2021) which are key tools for decoding intermediate token representations in transformer models. The logit lens applies the language modeling head to earlier layers without additional training, while the tuned lens in addition to this trains an affine mapping to align intermediate states with final token predictions. Direct logit attribution attributes logits to individual attention heads. This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs like Gemma-2, when prompted in non-latin language, processes via romanized latent states before producing non-English text. Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final non-English outputs, potentially masking the phenomenon under investigation.
Activation patching (Meng et al., 2022) is a key interpretability technique employed in our study. This technique has been used to draw causal interpretations of LLMs representations Variengien and Winsor (2023); Geiger et al. (2022); Kramár et al. (2024); Ghandeharioun et al. (2024); Chen et al. (2024). Building on these approaches, we adopt an activation patching-based experimental framework as defined in Dumas et al. (2024) to investigate and compare how concepts are encoded in romanized versus native scripts.
Previous studies have demonstrated that romanization can serve as an effective approach to interact with LLMs Husain et al. (2024). Liu et al. (2024) and Xhelili et al. (2024) employ an approach based on contrastive learning for post-training alignment, contrasting sentences with their transliterations in Latin script to overcome the script barrier and enhance cross-lingual transfer.
@@ 3 Background
Transformer’s forward pass: Here, we study autoregressive decoder-only transformer models. Vaswani (2017). Such models maintain the dimensionality of the input data throughout the computation process. For an input sequence , where is the sequence length, each token is initially mapped to an embedding vector, which is then transformed through the model’s layers, producing a sequence of hidden states (also referred to as latents).
The initial latents are obtained from a learned embedding matrix, which maps each vocabulary token to a fixed vector. These latents are updated iteratively across layers through the addition of residuals. The update rule for the latent at position in layer is expressed as:
The function , , is known as a transformer block, where is the final layer index. For prediction, the final embedding vector is multiplied by an unembedding matrix , producing a vector containing the logit scores for each vocabulary token . This is computed as:
These logit scores are converted to probabilities via the softmax function as follows:
@@ 4 Methodology
RQ1: Does LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
RQ2: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.2)
RQ3: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Latin languages? (Section 4.3)
4.1 Logit lens
Prompt design. In order to have the next token be easily inferred from the given prompt , we establish the following prompts.We use Hindi as the example language to describe the prompts. English translation and transliteration of the below prompts are present in Appendix A.
[…]
Repetition task. We ask the model to repeat the word mentioned in the same language.
[…]
Cloze task. We ask the model to predict the masked word in a sentence. We give the model two examples and ask it to predict the masked word in the next masked sentence.
[…]
Latent Romanization Analysis
Translation, repetition, and cloze tasks are explored by providing the respective prompts as inputs to an LLM to generate the corresponding output word. We romanize the output word, tokenize it, retaining only the tokens present in the model’s vocabulary, and analyze the occurrence of these tokens in the latent layers across timesteps of the output word generation. The analysis is done using logit lens by examining whether the probability of a romanized token in the next token distribution at a given layer exceeds 0.1 - a threshold optimized for detection accuracy i.e minimizing false positives and maximizing true positives. We refer to this hereafter as the latent romanization condition. Our analysis focuses on the final 10 layers of an LLM, where coherent romanized representations emerge according to logit lens visualizations (c.f. Figures 1 and 9 - 13).
We track romanized tokens using a timestep-specific tokenization scheme optimized for detection accuracy. In the first output generation timestep, we check for tokens that include the full romanized word and its prefixes. During intermediate timesteps,we check for all possible substrings of the romanized word in the latent layers. In the final output generation timestep, we probe the presence of only the full romanized word and its suffixes as potential tokens (c.f. Appendix B).
Latent romanization is analyzed under three distinct scenarios for the last 10 layers of an LLM:
Constrained Word Generation: Using standard prompts and the target word, we guide the model to generate the complete target word. At each layer, we track how often romanized tokens emerge during the decoding process. We do this by checking each generation timestep for latent romanization condition. The ‘latent fraction’ for a layer represents how frequently these romanized tokens appear across timesteps, averaged across all samples (c.f. Appendix C).
@@ 4.1 Logit lens
This analysis examines translations into target languages in native script and their romanized equivalents, using the translation prompts from Section 4.1. We focus on first-token generation of the output word, also considering possible synonyms.
In the next token generation step, the probability of target language and latent language (English) Wendler et al. (2024) at each layer is examined using logit lens. Each probability is computed by summing over probabilities of all possible tokens corresponding to the answer word(s) in that respective language (c.f. Appendix D). Tokens of latent language and target language are derived using tokenization scheme for first token generation timestep mentioned in Section 4.1.
4.3 Activation Patching
Prompt design. The prompt template used here is identical to the translation prompt employed for the logit lens (Section 4.1).
Approach: We intend to compare how concepts are encoded in native script versus romanized script. In order to do this, we patch representations from source prompt where input language is romanized and compare this with patching from source prompt where input language is in native script. We do this for patching with single source prompt and repeat it for patching with mean over multiple source prompt, where multiple romanized source input languages are contrasted with multiple native script source input languages. Single-source patching might be influenced by language or script-specific characteristics. In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language. We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f. Appendix E). We analyze to evaluate how effectively a concept is encoded in a given source input language .
@@ 5 Experimental Settings
5.1 Languages
We focus on five Indic languages: Hindi, Gujarati, Tamil, Telugu and Malayalam. Among these, Hindi and Gujarati belong to the Indo-Aryan branch of the Indo-European language family and use scripts derived from the Devanagari and Gujarati scripts, respectively. Tamil, Telugu, and Malayalam, on the other hand, are part of the Dravidian language family and each has its own distinct script. We also investigate latent romanization for Chinese and Georgian. Chinese belongs to the Sino-Tibetan language family and is written using logographic characters. Georgian is part of the Kartvelian language family and uses the unique Georgian script.
To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic. Greek is part of the Hellenic branch of the Indo-European language family and is written using the Greek alphabet. Ukrainian belongs to the East Slavic group of the Indo-European family and employs the Ukrainian alphabet, a variant of the Cyrillic script. Amharic is a South Ethio-Semitic language within the Afroasiatic family and is written using the Ge’ez script, an abugida where each character represents a consonant-vowel combination. Hebrew is a Northwest Semitic language within the Afroasiatic family and is written using the Hebrew alphabet, an abjad script originating from the Aramaic alphabet. Arabic is a Central Semitic language, also part of the Afroasiatic family, and utilizes the Arabic script, another abjad that evolved from the Nabataean alphabet. Notably, both Hebrew and Arabic scripts are written from right to left.
5.2 Language Models
In this study, we focus mainly on Gemma-2 9b, Gemma-2 9b instruction-tuned Team et al. (2024), Llama-2 7b and Llama-2 13b Touvron et al. (2023) language models, some of the best performing open weights English-centric LLMs. Although the training data for these models are primarily English, these models have high multilingual capabilities Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023).
| Native (input language in source prompt is in native script) | Romanized (input language in source prompt is in romanized script) | KL Divergence |
@@ 5.2 Language Models
5.3 Romanization
Transliteration is the conversion of text written in one script to another. Romanization is a subcategory of transliteration where the target script is English/Latin. We have taken into consideration two romanization schemes for Indic languages: (a) ITRANS scheme from IndicNLP library Kunchukuttan (2020) and (b) IndicXlit scheme Madhani et al. (2023). Based on our initial experiments, we observed that the IndicXlit scheme produces better romanization than ITRANS scheme. Hence for transliteration we have used the IndicXlit scheme Madhani et al. (2023). It generates romanization as is commonly used by native speakers and is trained on parallel transliteration corpora. We used pypinyin Mozillazg (2024) for Chinese and Unidecode Šolc (2025) for Georgian to romanize native scripts.
5.4 Logit Lens Experiments: Data
5.5 Activation Patching Experiments: Data
6 Results
@@ 6.1 Latent romanization
Our analysis demonstrates that LLMs do exhibit latent romanization during text completion tasks in six out of seven quantitatively analyzed languages, with Chinese being the only exception (c.f. 4) where this phenomenon is not observed. In Figure 4 we can see the latent fraction of romanized tokens for the last 10 layers of an LLM. This is across all tokens of the output word. The frequency of romanized tokens tends to increase just before the last layers. Qualitative logit lens analysis done for Greek, Ukrainian, Arabic, Amharic and Hebrew reveals similar patterns (c.f. Appendix F).
In Figure 5 it is observed that the range of the latent fraction of romanized tokens varies from in the first token generation step to in the last token generation step in most languages. This trend indicates that latent romanization increases progressively from the initial token to the final token of the output across languages. This observation supports our hypothesis that the first token generation involves more intricate decision-making processes compared to the generation of the final token within an output word.
@@
Among the three text completion tasks considered in the latent romanization experiment (Section 4.1), a relatively high latent romanization frequency is observed for the repetition task (Figure 6). The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic. This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers. This behavior could be attributed to the distribution of language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al. (2024).
As depicted in Figure 7, the probability distributions of source concepts in target language show remarkable similarity whether the source input language is in romanized or native script, consistent across both single-source and multi-source prompt setups. This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script. Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Latin languages in their romanized form as in their native script.
@@ 6.2 Patching with romanized representation versus native representation
We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens. In Figure 8, panels 8(b) and 8(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward. Conversely, in panels 8(a) and 8(c), where the target language is in romanized script, target tokens appear 1–2 layers earlier. This pattern indicates that when processing non-Latin languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script. This trend is consistent across language pairs and models, as shown in Figures 16-18. Additionally, native script curves (Figures 8(b), 8(d)) exhibit steeper gradients than their romanized equivalents (Figures 8(a), 8(c)).

### v2

# RomanLens: The Role Of Latent Romanization In Multilinguality In LLMs

@@ Abstract
Large Language Models (LLMs) exhibit remarkable multilingual generalization despite being predominantly trained on English-centric corpora. A fundamental question arises: how do LLMs achieve such robust multilingual capabilities? We take the case of non-Roman script languages, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a bridge in multilingual processing. Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in Romanized form before transitioning to native script, a phenomenon we term Latent Romanization. Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and Romanized scripts, suggesting a shared underlying representation. Additionally, for translation into non-Roman script languages, our findings reveal that when the target language is in Romanized form, its representations emerge earlier in the model’s layers compared to native script. These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of Romanization in facilitating language transfer.
RomanLens: The Role Of Latent Romanization In Multilinguality In LLMs
@@
Anoop Kunchukuttan1,6,
@@ 1 Introduction
In this study, we focus on languages written in non-Roman scripts. We hypothesize that LLMs leverage romanized forms of non-Roman script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations. Romanization—the representation of non-Roman scripts using Roman characters—may facilitate this process by aligning non-English languages more closely with English. Supporting this, Jaavid et al. (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations. We investigate whether LLMs indeed use romanization as a bridge between language-agnostic concepts and language-specific outputs given its potential implications for understanding multilingual processing in LLMs.
@@
To further probe this phenomenon, we employ activation patching Ghandeharioun et al. (2024); Variengien and Winsor (2024); Chen et al. (2024); Dumas et al. (2024), a technique that replaces activations from one forward pass with another to analyze the resulting outputs (c.f. Section 3.2). Dumas et al. (2024) found that LLMs process language and conceptual information as distinct entities. Building on this, we perform layerwise activation patching between romanized and native-script inputs to examine whether LLMs encode conceptual information similarly across scripts.
, ’pa’ is its romanized representation). Note that push is a prefix of pushpam, the romanized form of the translation of “flower”; also, poo is another romanized translation of “flower”.Color represents entropy of next-token generation from low (blue) to high (red). Plotting tool: Belrose et al. (2023).
@@
1. Latent Romanization Across Layers: During multilingual next-token generation, intermediate layers occasionally represent tokens in Romanized form before resolving to native script (Figure 1). We term this phenomenon Latent Romanization.
2. Consistent Semantic Encoding Across Scripts: Activation patching experiments reveal that LLMs encode semantic concepts similarly, regardless of whether the input is in native or Romanized script.
3. Earlier Emergence of Target Representations: When translating into Romanized versus native script, Romanized target representations emerge earlier in the model’s layers—typically one or two layers prior to native script representations (c.f. Figure 2).
@@ 2 Related Work
Recent studies have explored various aspects of LLMs’ multilingual behavior: examining whether English emerges as a latent language in English-centric LLMs Wendler et al. (2024), how the composition of training corpus mixtures influences latent representations Zhong et al. (2024) and how LLMs handle multilingual capabilitiesZhao et al. (2024b). Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al. (2023) and direct logit attribution Elhage et al. (2021) which are key tools for decoding intermediate token representations in transformer models. The logit lens applies the language modeling head to earlier layers without additional training, while the tuned lens in addition to this trains an affine mapping to align intermediate states with final token predictions. Direct logit attribution attributes logits to individual attention heads. This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman language, processes via romanized latent states before producing native language text. Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final native script outputs, potentially masking the phenomenon under investigation.
Activation patching (Meng et al., 2022) is a key interpretability technique employed in our study. This technique has been used to draw causal interpretations of LLMs representations Variengien and Winsor (2024); Geiger et al. (2022); Kramár et al. (2024); Ghandeharioun et al. (2024); Chen et al. (2024). Building on these approaches, we adopt an activation patching-based experimental framework to investigate and compare how concepts are encoded in romanized versus native scripts.
Previous studies have demonstrated that romanization can serve as an effective approach to interact with LLMs Jaavid et al. (2024). Liu et al. (2024) and Xhelili et al. (2024) employ an approach based on contrastive learning for post-training alignment, contrasting sentences with their transliterations in Roman script to overcome the script barrier and enhance cross-lingual transfer.
@@ 3 Background
Transformer’s Forward Pass
Decoder-only transformer models (Vaswani, 2017) employ a residual architecture to process input sequences through multiple layers, producing a sequence of hidden states (latents). These latents, whose dimensionality remains the same are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index. For next-token prediction, the final latent is transformed by an unembedding matrix to produce logit scores for vocabulary tokens which are then converted to probabilities via the softmax function (c.f. Appendix A).
@@ 4 Methodology
RQ1: Do LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman languages? (Section 4.2)
RQ3: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.3)
Prompt design. We design prompts that facilitate next-token () prediction from the given context (). This is adopted across all analysis setups. Appendix B provides a Hindi example of this prompt design, including its English translation and romanized form.
Translation task. We prompt the model to translate a word given five in-context examples.
Repetition task. We prompt the model to repeat a word in the same language given 5 in-context examples.
Cloze task. We prompt the model to predict the masked word in a sentence given two in-context examples.
4.1 Latent Romanization Analysis
Translation, repetition, and cloze tasks are explored by providing the respective prompts as inputs to an LLM to generate the corresponding output word. We romanize the output word, tokenize it, retaining only the tokens present in the model’s vocabulary, and analyze the occurrence of these tokens in the latent layers across timesteps of the output word generation. The analysis is done using logit lens by examining whether the probability of a romanized token in the next token distribution at a given layer exceeds 0.1. We refer to this hereafter as the latent romanization condition. The 0.1 threshold is empirically determined to optimize detection accuracy, i.e. minimizing false positives and maximizing true positives (compared to alternative thresholds 0.05 and 0.01). Our analysis focuses on the final 10 layers of an LLM, where coherent romanized representations emerge according to logit lens visualizations (c.f. Figures 1 and 8 - 12).
We track romanized tokens using a timestep-specific tokenization scheme optimized for detection accuracy. In the first output generation timestep, we check for tokens that include the full romanized word and its prefixes. During intermediate timesteps,we check for all possible substrings of the romanized word in the latent layers. In the final output generation timestep, we probe the presence of only the full romanized word and its suffixes as potential tokens (c.f. Appendix C).
Latent romanization is analyzed under three distinct scenarios:
Constrained Word Generation: Using standard prompts and the target word, we guide the model to generate the complete target word. At each layer, we track how often romanized tokens emerge during the decoding process. We do this by checking each generation timestep for latent romanization condition. The ‘latent fraction’ for a layer represents how frequently these romanized tokens appear across timesteps, averaged across all samples (c.f. Appendix D).
@@ 4.2 Patching With Romanized Representation Versus Native Representation
We intend to compare how concepts are encoded in native script versus romanized script using translation task. In order to do this, we patch representations from source prompt where input language is romanized and compare this with patching from source prompt where input language is in native script. We first perform patching using a single-source prompt. Then, we repeat the process using an averaged multi-source prompt, contrasting multiple romanized source input languages with multiple native script source input languages. Single-source patching might be influenced by language or script-specific characteristics. In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language. We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f. Appendix G). We analyze to evaluate how effectively a concept is encoded in a given source input language .
@@
This analysis examines translation task with target languages in native script and their romanized equivalents. We focus on first-token generation of the output word, also considering possible synonyms.
In the next token generation step, the probability of target language and latent language (English) Wendler et al. (2024) at each layer is examined using logit lens. Each probability is computed by summing over probabilities of all possible tokens corresponding to the answer word(s) in that respective language (c.f. Appendix E). Tokens of latent language and target language are derived using tokenization scheme for first token generation timestep mentioned in Section 4.1.
@@ 5 Experimental Settings
Languages:
We focus on five Indic languages: Hindi, Gujarati, Tamil, Telugu, and Malayalam, as well as Chinese and Georgian. Among these, Hindi and Gujarati belong to the Indo-Aryan branch of the Indo-European language family and use scripts derived from the Devanagari and Gujarati scripts, respectively. Tamil, Telugu, and Malayalam, on the other hand, are part of the Dravidian language family and each has its own distinct script. Chinese belongs to the Sino-Tibetan language family and is written using logographic characters. Georgian is part of the Kartvelian language family and uses the unique Georgian script. To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic (c.f Appendix H).
Language Models:
In this study, we focus mainly on Gemma-2 9b, Gemma-2 9b instruction-tuned Team et al. (2024), Llama-2 7b, Llama-2 13b Touvron et al. (2023) and Mistral- 7B Jiang et al. (2023) (c.f. Appendix K) language models, some of the best performing open weights English-centric LLMs. Although the training data for these models are primarily English, these models have high multilingual capabilities Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023).
Romanization:
Data For Logit Lens Experiments:
Data For Activation Patching Experiments:
6 Results
We now share our observations about the role of Romanization in LLMs.
| Native | Romanized | |
@@ 6.1 Latent romanization
Our analysis demonstrates that LLMs do exhibit latent romanization during text completion tasks in six out of seven quantitatively analyzed languages (c.f. Figure 4(a)), with Chinese being the only exception where this phenomenon is not observed. In Figure 4(a) we can see the latent fraction of romanized tokens for the last 10 layers of an LLM. This is across all tokens of the output word. The frequency of romanized tokens tends to increase just before the last layers. Qualitative logit lens analysis done for Greek, Ukrainian, Arabic, Amharic and Hebrew reveals similar patterns (c.f. Appendix H).
In Figures 4(b) and 4(c) it is observed that the range of the latent fraction of romanized tokens varies from in the first token generation step to in the last token generation step in most languages. This trend indicates that latent romanization increases progressively from the initial token to the final token of the output across languages. This observation supports our hypothesis that the first token generation involves more intricate decision-making processes compared to the generation of the final token within an output word.
@@
Among the three text completion tasks considered in the latent romanization experiment (Section 4.1), a relatively high latent romanization frequency is observed for the repetition task (Figure 6). The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic. This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers. This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al. (2024).
6.2 Patching With Romanized Representation vs. Native Representation
In Figure 5, we analyze two patching scenarios: In the single-source setup, we compare patching from Hindi→Italian with Hindi(romanized)→Italian source prompt, to Malayalam→Italian target prompt. In the multi-source setup, we contrast patching from multiple native script prompts (Hindi→Italian, Gujarati→Italian ) against their romanized counterparts (Hindi(romanized)→Italian, Gujarati(romanized)→Italian ). We compare the probability distributions of source concept in target language across adjacent graphs where native source input language is contrasted with romanized source input language. It is evident that these probability distributions show remarkable similarity whether the source input language is in romanized or native script, consistent across both single-source and multi-source prompt setups. The similarity is quantitatively supported by the KL divergence measurements between adjacent graphs, remaining below 0.01 in both setups. KL divergence value close to zero indicates that the two distributions are nearly identical.
This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script. Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman languages in their romanized form as in their native script.
@@
We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens. In Figure 7, panels 7(c) and 7(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward. Conversely, in panels 7(a) and 7(b), where the target language is in romanized script, target tokens appear 1–2 layers earlier. This pattern indicates that when processing non-Roman script languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script. This trend is consistent across language pairs and models, as shown in Figures 16-18. This suggests that romanization facilitates faster progression toward language-specific embeddings.
In Figure 7, in all four graphs consistent with Wendler et al. (2024), English representations emerge from the middle layers and persist until the final few layers, where the target language representations gradually take shape. It is important to note that native script curves (Figures 7(c), 7(d)) exhibit steeper gradients than their romanized equivalents (Figures 7(a), 7(b)).
Discussion.
In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks. Based on previous works in this field Wendler et al. (2024); Zhao et al. (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language. Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Roman script languages, an observation strongly supported by our analysis of six diverse writing systems. Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al. (2024); Xhelili et al. (2024) work. Notably, we do not observe Latent Romanization in Chinese, likely due to its logographic script and relatively high-resource status.